# A Unified Framework for 3D Scene Understanding

### v1

@@ 4.2 Analysis and Ablation
We conduct ablation studies and analyze the key insights of our designs. All models are evaluated on unified tasks to better show the effectiveness of the proposed components on wide scopes.
@@
Design of inter-task associations. In our approach, we use knowledge distillation and contrastive learning to connect supported tasks. As shown in Tab. 2, when applying the distillation, i.e. row $2$, the performance of IS and interactive segmentation increase to $59.3$ mAP and $56.6$ AP, respectively. We believe the improvement on the IS is because of the reliable knowledge distilled from the interactive segmentation, and the improvement on interactive segmentation is attributed to the intrinsic connections between the two tasks. Then, we ablate the ranking-based contrastive learning, i.e. row $3$. We observe improvements on four tasks, including the generic segmentation and the referring segmentation, while a bit of performance drop on the interactive segmentation. This phenomenon suggests that the contrastive learning is effective in most tasks while there is a huge struggle to align the point and text modalities, which weakens the interactive segmentation performance. Applying the knowledge distillation and the ranking-based contrastive learning simultaneously obtains comparable performance on six tasks, indicating the complementarity of the two components. We further employ the two-stage fine-tuning trick, bringing consistent improvements across various tasks.
@@
Detailed ablation on the components is shown in Tab. 3. It is observed that knowledge distillation to various tasks brings respective improvements. As for the contrastive learning, comparing row $2$ and row $4$ in Tab. 3(b), the ranking rule suppresses the confusing point-text pairs, boosting the contrastive learning to be more effective. $\lambda$ controls the strength of the explicit inter-task associations. We empirically find that setting $\lambda$ to 0.1 obtains the best performance, as shown in Tab. 5.
Influence of task unification. The proposed framework is flexible and can be conveniently extended to additional tasks. As shown in Tab. 5, we ablate the number of unified tasks and observe a continuous performance decline on the PS, IS, and interactive segmentation. It demonstrates a significant challenge in balancing different tasks.
@@
Influence of vision prompts. We empirically find that the vision prompts affect the interactive segmentation performance. For a fair comparison, we use the same vision prompts generated by the AGILE3D yue2023agile3d to evaluate our interactive segmentation performance.
Furthermore, to analysis the influence of vision prompts, we ablate the 3D spatial distances between the vision prompts and the instance centers. Specifically, assuming an instance containing $n$ points, we denote the mean coordinate of these points as the instance center and order the $n$ points based on their distances to the instance center. Then, we evaluate the interactive segmentation performance while using the $\left\lfloor D\times n\right\rfloor$-th nearest point as the vision prompt, as shown in Tab. 4. When the vision prompt is located at the instance center, the interactive segmentation achieves the upper bound performance of $56.6$ AP. There is a significant performance gap (up to $20.2$ AP) between the edge point and center point. It illustrates considerable room for improvement. We observe an unusual decline in AP while converting $D$ from $0.9$ to $1.0$. We think this is because of the ambiguity to distinguish the edge points from adjacent instances. As we all know, this is the first work ablating the influence. We will explore it deeply in the future work.
@@ 5 Conclusion and Discussion
We design a unified framework named UniSeg3D, which offers a flexible and efficient solution for 3D scene understanding, supporting six tasks within a single model. Previous task-specific approaches fail to extract across-task information, limiting their understanding of 3D scenes to a task-specific perspective. In contrast, we take advantage of the multi-task design and enhance the performance through building inter-task associations. Specifically, we propose the knowledge distillation and ranking-based contrastive to facilitate across-task knowledge sharing. Experiments demonstrate that the proposed framework is powerful and achieves SOTA performance on all six unified tasks.
Limitation. The proposed method aims to achieve unified 3D scene understanding. However, this work focuses on indoor tasks, and lacks discussions in outdoor scenes. The modality gap between the 3D point cloud and the linguistic texts is still significant, leading to performance discrepancy between the 3D interactive segmentation and the 3D referring segmentation. For the interactive segmentation task, we find out that UniSeg3D performs worse when the vision prompt is located away from the instance centers. It limits the reliability of the UniSeg3D and should be explored in the future work.
@@
  • Appendix B: Analysis on the inference time of the proposed UniSeg3D.
  • Appendix C: Qualitative visualizations in more 3D scenarios.
A Comparisons with Approaches Specialized for Specific Tasks.
The experiments presented in the main manuscript primarily utilize the overall metrics each task. This section provides a comprehensive comparison of each task using more detailed metrics. The performance of 3D open-vocabulary semantic segmentation is assessed on the ScanNet200 by utilizing masks generated from UniSeg3D  trained on ScanNet20. Followingnguyen2023open3dis , 51 classes that are semantically close to the ScanNet20 categories are grouped as “Base”, while the remaining classes are grouped as “Novel”.
Figure I: For comparison with existing instance segmentation methods, we test our approach on ScanNet, achieving performance that is competitive with state-of-the-art.
Method Reference Instance
mAP25 mAP50 mAP
3D-SIShou20193dsis CVPR’19 35.7 18.7 -
GSPNyi2019gspn CVPR’19 53.4 37.8 19.3
PointGroupjiang2020pointgroup CVPR’20 71.3 56.7 34.8
OccuSeghan2020occuseg CVPR’20 71.9 60.7 44.2
DyCo3Dhe2021dyco3d CVPR’21 72.9 57.6 35.4
SSTNetliang2021sstnet ICCV’21 74.0 64.3 49.4
HAISchen2021hais ICCV’21 75.6 64.4 43.5
DKNetwu2022dknet ICCV’22 76.9 66.7 50.8
SoftGroupvu2022softgroup CVPR’22 78.9 67.6 45.8
PBNetzhao2023pbnet ICCV’23 78.9 70.5 54.3
ISBNetngo2023isbnet CVPR’23 82.5 73.1 54.5
SPFormersun2023superpoint AAAI’23 82.9 73.9 56.3
Mask3Dschult2023mask3d ICRA’23 83.5 73.7 55.2
MAFTlai2023mask ICCV’23 - 75.9 58.4
QueryFormerlu2023query ICCV’23 83.3 74.2 56.5
OneFormer3Dkolodiazhnyi2023oneformer3d CVPR’24 86.4 78.1 59.3
@@
Figure II: Comparison with previous 3D open-vocabulary semantic segmentation methods on ScanNe200. Experimental results demonstrate that our work outperforms the state-of-the-art approaches in terms of AP.
@@
OpenScene peng2023openscene + Mask3D schult2023mask3d CVPR’23 7.6 11.1 8.5
OpenMask3Dtakmaz2024openmask3d NeurIPS’23 11.9 14.3 12.6
SOLElee2024segment CVPR’24 19.1 17.4 18.7
Open3DISnguyen2023open3dis CVPR’24 16.5 25.8 19.0
UniSeg3D (ours) - 18.0 24.4 19.7
Figure III: Comparison with previous 3D interactive segmentation methods on ScanNet. UniSeg3D demonstrates remarkable performance in AP, AP50 and AP25.
@@
InterObject3D kontogianni2023interactive ICRA’23 20.9 38.0 67.2
AGILE3D yue2023agile3d ICLR’24 53.5 75.6 91.3
UniSeg3D (ours) - 54.5 79.4 93.2
Figure IV: Comparison with previous 3D referring segmentation methods on ScanRefer. UniSeg3D demonstrates significant performance in terms of mIoU and acc@0.25
Method Reference mIoU acc@0.5 acc@0.25
TGNN huang2021text AAAI’21 24.9 28.2 33.2
X-RefSeg3D qian2024x AAAI’24 25.5 28.6 34.0
UniSeg3D (ours) - 29.6 28.0 41.5
B Analysis on Inference Time.
This work proposes a unified framework, achieving six tasks in once inference. It is more efficient compared with running six approaches specialized for a single task. We present the inference time of the proposed method for further analysis of efficiency. Tab. I illustrates that our method achieves effective integration across six tasks while maintaining highly competitive inference times compared to previous methods.
Table I: The inference time and instance segmentation accuracy on the ScanNet validation split.
@@
C Qualitative Visualizations.
We provide additional qualitative results in this section. In Fig. V, visualizations of segmentation results for additional scenes are presented, showcasing the point clouds, the ground truth, and the segmentation result obtained by UniSeg3D for all $6$ tasks within each scene. In Fig. VI, we present the visualizations of segmentation results obtained by UniSeg3D and the current SOTA methods on the ScanNet dataset. We list four current methods, including OneFormer3D, AGILE3D, X-RefSeg3D, and Open3DIS, each of which achieved SOTA exclusively within their respective tasks, while UniSeg3D accomplished high-quality segmentation across all tasks.
Figure V: Visualization of segmentation results obtained by UniSeg3D on the ScanNet validation split.
Figure VI: Visualization of segmentation results obtained by UniSeg3D and current SOTA methods on the ScanNet validation split.

### v2

@@ 4.2 Analysis and Ablation
We conduct ablation studies and analyze the key insights of our designs. All models are evaluated on unified tasks to show the effectiveness of the proposed components on a broad scope.
The challenge of multi-task unification. We first discuss the challenge of unifying multi-tasks in a single model. Specifically, we simply add interactive segmentation, referring segmentation, and OVS into our framework to build a unification baseline, as shown in Tab. 2. We observe a continuous performance decline on the PS, IS, and interactive segmentation tasks, indicating a significant challenge in balancing different tasks. Even so, we believe that unifying multiple tasks within a single model is worthy of exploring, as it can reduce computation consumption and benefit real-world applications. Thus, this paper proposes to eliminate performance decline by delivering inter-task associations, and the following experiments demonstrate that this could be a valuable step.
Table 3: Ablation on components. “Distillation”, “Rank-Contrastive”, and “Trick” denote the knowledge distillation, ranking-based contrastive learning, and two-stage fine-tuning trick, respectively.
Datasets ScanNet20 ScanRefer ScanNet200 Overall
Components Pan. Sem. Inst. Inter. Ref. OV
Distillation Rank-Contrastive Trick PQ mIoU mAP AP mIoU AP
- - - 70.4 76.2 58.0 54.5 29.1 19.7 51.3
- - 70.9 76.2 58.6 55.3 29.2 19.6 51.6
- - 70.8 76.4 58.4 54.1 29.6 19.9 51.5
- 71.3 76.3 59.1 54.1 29.5 19.6 51.7
71.3 76.9 59.3 54.5 29.6 19.7 51.9
Table 4: Ablation on different designs of the proposed components. “${v\to g}$” and “${v\to r}$” denote the knowledge distillation from the interactive segmentation to the generic segmentation and the referring segmentation, respectively. “Contrastive” and “Rank” denote the contrastive learning and the ranking rule, respectively.
(a) Ablation on designs for knowledge distillation.
Datasets ScanNet20 ScanRefer ScanNet200 Overall
Components Pan. Sem. Inst. Inter. Ref. OV
${v\to g}$ ${v\to r}$ PQ mIoU mAP AP mIoU AP
- - 70.8 76.4 58.4 54.1 29.6 19.9 51.5
- 71.2 76.3 59.0 54.0 29.5 19.8 51.6
- 70.7 76.2 58.6 54.1 29.7 20.0 51.6
71.3 76.3 59.1 54.1 29.5 19.6 51.7
(b) Ablation on designs for ranking-based contrastive learning.
Datasets ScanNet20 ScanRefer ScanNet200 Overall
Components Pan. Sem. Inst. Inter. Ref. OV
Contrastive Rank PQ mIoU mAP AP mIoU AP
- - 70.9 76.2 58.6 55.3 29.2 19.6 51.6
- 71.0 76.3 59.0 54.5 29.4 19.7 51.7
- 71.0 76.2 58.7 54.6 29.5 19.8 51.6
71.3 76.3 59.1 54.1 29.5 19.6 51.7
@@
Design of inter-task associations. Our approach uses knowledge distillation and contrastive learning to connect supported tasks. As shown in Tab. 3, when applying the distillation, i.e. row $2$, the performance of IS and interactive segmentation increase to $58.6$ mAP and $55.3$ AP, respectively. We believe the improvement on the IS task is because of the reliable knowledge distilled from the interactive segmentation, and the improvement on the interactive segmentation task is attributed to the intrinsic connections between the two tasks. Then, we ablate the ranking-based contrastive learning, i.e. row $3$. We observe improvements on five tasks, including the generic segmentation and the referring segmentation, while a bit of performance drop on the interactive segmentation. This phenomenon suggests that contrastive learning is effective in most tasks, but there is a huge struggle to align the point and text modalities, which weakens the interactive segmentation performance. Overall metric measures multi-task unification performance. We choose models and checkpoints with higher Overalls in our experiments. In practical applications, checkpoints can be chosen based on preferred tasks while maintaining good performance across other tasks. Applying knowledge distillation and ranking-based contrastive learning obtains comparable performance on most tasks, performing higher Overall than rows $2$ and $3$, indicating the complementarity of the two components. We further employ two-stage fine-tuning trick, bringing consistent improvements across various tasks.
Table 5: Ablation on hyper-parameter $\lambda$.
Datasets ScanNet20 ScanRefer ScanNet200 Overall
Hyper-parameter Pan. Sem. Inst. Inter. Ref. OV
$\lambda$ PQ mIoU mAP AP mIoU AP
0.05 70.7 76.2 58.9 54.4 29.5 19.6 51.6
0.1 71.3 76.3 59.1 54.1 29.5 19.6 51.7
0.2 70.8 76.6 58.6 52.3 29.8 19.5 51.3
0.3 70.6 75.7 58.4 51.6 29.6 19.3 50.9
@@
Detailed ablation on the components is shown in Tab. 4. It is observed that knowledge distillation to various tasks brings respective improvements. As for contrastive learning, comparing row $2$ and row $4$ in Tab. 4(b), the ranking rule suppresses the confusing point-text pairs, boosting contrastive learning to be more effective. $\lambda$ controls the strength of the explicit inter-task associations. We empirically find that setting $\lambda$ to $0.1$ obtains the best performance, as shown in Tab. 5.
@@
Influence of vision prompts. We empirically find that the vision prompts affect the interactive segmentation performance. To ensure a fair comparison, we adopt the same vision prompts generation strategy designed in AGILE3D yue2023agile3d to evaluate our interactive segmentation performance.
Furthermore, to analyze the influence of vision prompts, we ablate the 3D spatial distances between the vision prompts and the instance centers. Specifically, assuming an instance containing $n$ points, we denote the mean coordinate of these points as the instance center and order the $n$ points based on their distances to the instance center. Then, we evaluate the interactive segmentation performance while using the $\left\lfloor r_{d}\times n\right\rfloor$-th nearest point as the vision prompt, as shown in Tab. 6. When the vision prompt is located at the instance center, the interactive segmentation achieves the upper bound performance of $56.6$ AP. There is a significant performance gap (up to $20.2$ AP) between the edge and center points. It illustrates considerable room for improvement. We observe an unusual decline in AP while increasing $r_{d}$ from $0.9$ to $1.0$. We think this is because of the ambiguity in distinguishing the edge points from adjacent instances. As we all know, this is the first work ablating the influence of vision prompts. We will explore it in depth in future work.
@@ 5 Conclusion and Discussion
We propose a unified framework named UniSeg3D, which provides a flexible and efficient solution for 3D scene understanding, supporting six tasks within a single model. Previous task-specific approaches fail to leverage cross-task information, limiting their understanding of 3D scenes to a task-specific perspective. In contrast, we take advantage of the multi-task design and enhance performance through building inter-task associations. Specifically, we employ knowledge distillation and ranking-based contrastive learning to facilitate cross-task knowledge sharing. Experiments demonstrate the proposed framework is a powerful method, achieving SOTA performance across six unified tasks.
Limitation. UniSeg3D aims to achieve unified 3D scene understanding. However, it works on indoor tasks and lacks explorations in outdoor scenes. Additionally, we observe that UniSeg3D performs worse interactive segmentation performance when the vision prompt is located away from the instance centers, limiting the reliability of the UniSeg3D and should be explored in future work.
@@
  • Appendix B: Inference time analysis of the proposed UniSeg3D.
  • Appendix C: Qualitative visualizations illustrating model effectiveness.
Appendix A Comparisons employing more metrics on specific tasks.
The experiments presented in the main manuscript primarily use overarching metrics to measure performance on each task. This section provides more comprehensive comparisons of our method on each task using detailed metrics. We train the model on ScanNet20 and assess its open-vocabulary segmentation performance on ScanNet200. Following nguyen2023open3dis , $51$ classes in ScanNet200 that are semantically similar to annotated classes in ScanNet20 are grouped as Base classes, while the remaining classes are divided as Novel classes. The model is then directly tested on Replica straub2019replica to evaluate its zero-shot segmentation performance.
Figure I: Comparison with existing instance segmentation methods on ScanNet20. UniSeg3D achieves highly competitive performance.
Method Reference mAP25 mAP50 mAP
3D-SIShou20193dsis CVPR 19 35.7 18.7 -
GSPNyi2019gspn CVPR 19 53.4 37.8 19.3
PointGroupjiang2020pointgroup CVPR 20 71.3 56.7 34.8
OccuSeghan2020occuseg CVPR 20 71.9 60.7 44.2
DyCo3Dhe2021dyco3d CVPR 21 72.9 57.6 35.4
SSTNetliang2021sstnet ICCV 21 74.0 64.3 49.4
HAISchen2021hais ICCV 21 75.6 64.4 43.5
DKNetwu2022dknet ICCV 22 76.9 66.7 50.8
SoftGroupvu2022softgroup CVPR 22 78.9 67.6 45.8
PBNetzhao2023pbnet ICCV 23 78.9 70.5 54.3
ISBNetngo2023isbnet CVPR 23 82.5 73.1 54.5
SPFormersun2023superpoint AAAI 23 82.9 73.9 56.3
Mask3Dschult2023mask3d ICRA 23 83.5 73.7 55.2
MAFTlai2023mask ICCV 23 - 75.9 58.4
QueryFormerlu2023query ICCV 23 83.3 74.2 56.5
OneFormer3Dkolodiazhnyi2023oneformer3d CVPR 24 86.4 78.1 59.3
@@
Figure II: Comparison with previous 3D interactive segmentation methods on ScanNet20. UniSeg3D presents remarkable performance in terms of three metrics.
@@
InterObject3D kontogianni2023interactive ICRA 23 20.9 38.0 67.2
AGILE3D yue2023agile3d ICLR 24 53.5 75.6 91.3
UniSeg3D (ours) - 54.5 79.4 93.2
Figure III: Comparison with existing 3D referring segmentation methods on ScanRefer. UniSeg3D demonstrates notable performance in terms of mIoU and acc@0.25.
Method Reference mIoU acc@0.5 acc@0.25
TGNN huang2021text AAAI 21 24.9 28.2 33.2
X-RefSeg3D qian2024x AAAI 24 25.5 28.6 34.0
UniSeg3D (ours) - 29.6 28.0 41.5
Table I: Comparison with previous open-vocabulary segmentation methods on ScanNe200 and Replica. Our method outperforms existing approaches in terms of AP.
Method Reference ScanNet200 Replica
@@
OpenScene peng2023openscene with schult2023mask3d CVPR 23 8.5 11.1 7.6 10.9 15.6 17.3
OpenMask3Dtakmaz2024openmask3d NeurIPS 23 12.6 14.3 11.9 13.1 18.4 24.2
SOLElee2024segment CVPR 24 18.7 17.4 19.1 - - -
Open3DISnguyen2023open3dis CVPR 24 19.0 25.8 16.5 18.5 24.5 28.2
UniSeg3D (ours) - 19.7 24.4 18.0 19.1 24.1 29.2
Appendix B Inference time analysis of the proposed UniSeg3D.
This work proposes a unified framework, achieving six tasks in one inference, which would be more efficient than running six task-specific approaches individually. We present the inference time of the proposed method for efficiency analysis. Tab. II illustrates that our method achieves effective integration across six tasks while maintaining highly competitive inference times compared to previous methods.
Table II: Inference time and instance segmentation performance on the ScanNet20 validation split.
@@
Appendix C Qualitative visualizations illustrating model effectiveness.
We provide qualitative results in this section. In Fig. IV, visualizations of multi-task segmentation results are presented, showcasing point clouds, ground truth, and predictions within each scene. In Fig. V, we present visualizations of predictions from UniSeg3D and current SOTA methods. In Fig. VI, we test our model on open-set classes not included in training data to evaluate the model’s open capability. Furthermore, we even replace the class names with attribute descriptions in the open vocabulary, and impressively, we observe the preliminary reasoning capabilities of our approach.
Figure IV: Visualization of segmentation results obtained by UniSeg3D on ScanNet20 validation split.
Figure V: Visualization of segmentation results obtained by UniSeg3D and current SOTA methods on ScanNet20 validation split.
Figure VI: Visualization of open capabilities. Red prompts involve categories not presented in the ScanNet200 annotations, while blue prompts describe the attributes of various objects, such as affordances and color.